6.   A. If the MPC is = to .9, what is the multiplier?  

        B. If the MPC is = to .75, what is the multiplier?

        C. If the MPC is = to .6, what is the multiplier?

        D. If the MPC is = to .5, what is the multiplier? 20 points

  1. If the MPC is = to .9, and spending increases by $30 billion, what is the increase in GDP? Show all work. 5 points
  2. If the MPC is = to .8, and spending increases by $30 billion, what is the increase in GDP? Show all work. 5 points

  1. If the MPC is = to .75, and spending increases by $30 billion, what is the increase in GDP? Show all work. 5 points
  1. If the MPC is = to .5, and spending increases by $30 billion, what is the increase in GDP? Show all work. 5 points
  1. Notice that in questions 7, 8, & 9 the MPC decreases, the increase in spending stays the same and the increase in GDP decreases. Why does that occur? 5 points
  1. If the MPC is = .9 and spending decreases by $100 billion, what happens to GDP? Show all work. 5 points.
Answer #1

6) The multiplier is calculated as: 1/1-MPC

a) When the MPC is .9

Multiplier = 1/1-.9

=1/.1

=10

b) When the MPC is .75

Multiplier = 1/1-.75

=1/.25

=4

c) When the MPC is .6

Multiplier = 1/1-.6

=1/.4

=2.5

d) When the MPC is .5

Multiplier = 1/1-.5

=1/.5

=2

Please upload the next question in a seprate post as per Cheg guidelines.Thanks